ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 122, 2001

Alon, N. and P. Pudlák Constructive lower bounds for off-diagonal Ramsey numbers

243

Beidar, K. I. and M. Brešar Extended Jacobson density theorem for rings with derivations and automorphisms

317

Brešar, M. See K. I. Beidar

Chen, Z. G. Estimates on μ(z)-homeomorphisms of the unit disk  347

Chernov, N. and D. Kleinbock Dynamical Borel–Cantelli lemmas for Gibbs measures

1

Dolfi, S. and M. S. Lucido Finite groups in which p′-classes have

q′-length 93

Efrat, I. A Hasse principle for function fields over PAC fields  43
Flicker, Y. Z. Motivic torsors  61

Giambruno, A. and P. Koshlukov On the identities of the Grassmann algebras in characteristic p > 0

305

Ginzburg, V. L. and A. Golubev Holonomy on Poisson manifolds and the modular class

221

Golubev, A. See V. L. Ginzburg

González, M. and A. Martínez-Abejón Ultrapowers of L_{1}(μ) and the subsequence splitting principle

189

Goren, E. Z. Hasse invariants for Hilbert modular varieties  157

Herron, D. and D. Minda Comparing invariant distances and conformal metrics on Riemann surfaces

207

Jordan, D. and M. Steiner Compact surfaces as configuration spaces of mechanical linkages

175

Kapovich, I. and D. T. Wise On the failure of the co-Hopf property for subgroups of word-hyperbolic groups

125

Kleinbock, D. See N. Chernov

Koshlukov, P. See A. Giambruno

Lindenstrauss, E. p-Adic foliation and equidistribution  29

Lucido, M. S. See S. Dolfi

Martínez-Abejón, A. See M. González

Minda, D. See D. Herron Mohnke, K. Lagrangian embeddings in the complement of symplectic

hypersurfaces

117

Naor, A. and A. Zvavitch Isomorphic embedding of ℓ_{p}^{n}, 1 < p < 2,

into ℓ_{1}^{(1+ϵ)n} 371

Puchta, J.-C. Groups with multiplicative subgroup growth  149

Pudlák, P. See N. Alon

Ran, Z. On the variety of rational space curves  359

Roessler, D. Lambda-structure on Grothendieck groups of hermitian vector bundles

279

Sajith, G. and S. Thangavelu On the injectivity of twisted spherical means on ℂ^{n}

79

Steiner, M. See D. Jordan

Thangavelu, S. See G. Sajith

Vershynin, R. John’s decompositions: Selecting a large part  253

Wise, D. T. See I. Kapovich

Zvavitch, A. See A. Naor